OpenStudy (anonymous):

In the function f(x) = 3x - 6 how does an increase of 1 in the independent variable x affect f(x)? f(x) will have an increase of 3 f(x) will have an increase of 9 f(x) will have a decrease of 3x - 3 f(x) will have a decrease of 3

OpenStudy (anonymous):

substitute 1 instead of the variable x: f(x)=3(1)-6=3-6=-3

OpenStudy (anonymous):

hence f(x) will have a decrease of 3